# Emergent classical gravity as stress-tensor deformed field theories

### v1

@@ Abstract
We introduce a framework in which classical gravity arises from stress-tensor deformations of quantum field theories (QFTs). By expressing the deformed partition function as a gravitational path integral, we propose an equivalent description between a seed QFT coupled to gravity and an effective field theory defined on a deformation-induced metric. This approach, which does not rely on holography or extra dimensions, applies to a variety of gravitational models, including Einstein and Palatini gravity. The deformation parameter acts as a coupling to the emergent geometry, linking stress-energy dynamics to spacetime curvature. The framework provides a calculable, dimension-independent method for studying gravitational theories resulting from QFT deformations.
@@ I Introduction
A longstanding question is whether classical gravity can emerge from microscopic quantum degrees of freedom [1]. Several lines of evidence point in this direction. Thermodynamic and entanglement arguments relate Einstein’s equations to equilibrium or near–equilibrium properties of quantum matter [2, 3]. Holographic dualities, most notably AdS/CFT [4, 5], provide concrete realizations of spacetime from large-$N$ QFT data, but rely on specific boundary conditions and asymptotics. Non-holographic approaches that reconstruct geometry directly from QFT therefore offer a broader setting [6, 7].
A particularly promising clue comes from deformations generated by the QFT stress tensor. In two dimensions, the $T\bar{T}$ deformation [8, 9] leads to a solvable flow and admits an interpretation as coupling to (nearly) topological gravity or gravitational dressing [10, 11, 12, 13] or random geometry [14]. These results show explicitly that the deformation of QFT by stress-tensor operators can be rephrased as a geometric response. Higher-dimensional stress tensor deformations have been proposed in [15, 16, 17], and the generic geometric realization has been demonstrated in [18]. However, beyond d=2d{=}2 and outside special kinematic regimes, there is no general, non-holographic mechanism that derives classical gravitational field equations from a purely field-theoretic construction. Related thought experiments and recent proposals suggest that matter dynamics alone may trigger flows toward gravitational behavior [19], but a calculable, dimension-independent framework remains to be formulated.
We propose a framework for studying stress-tensor–induced deformations of a seed QFT, where the deformed partition function is expressed as a path integral over metrics weighted by a gravitational action. This formulation leads to non-local stress tensor deformations of the field theory, which can be identified with local deformations of an effective action defined on a deformed metric, allowing us to regard the effective field action with the deformation as a corresponding gravitational action. The framework extends naturally to Palatini gravity [20, 21], suggesting that a broad class of classical gravitational theories can emerge from stress-tensor–driven deformations of QFTs.
@@ II Assumptions and Framework
The framework is based on relating a deformed QFT to a gravitational theory through the path integral. The gravitational partition function is defined as
@@
where $\hat{S}[g,\psi]$ denotes the seed theory action and $\psi$ collectively denotes the matter fields.
The metric integral is evaluated in the saddle-point approximation, giving
@@
$\sim\int\mathcal{D}\psi\sum_{\hat{\gamma}}\Big{(}\sum_{\alpha}\mathcal{Z}^{(\lambda)}_{\mathrm{loops}}[g^{*}_{\alpha}]\;e^{-S_{\alpha}^{(\lambda)}[\hat{\gamma},\psi]}\Big{)},$ (2)
where $\{g^{*}_{\alpha}\}$ denotes the set of all saddle points satisfying $\lim_{\lambda\to 0}g^{*}_{\alpha}=\hat{\gamma}$, and $\hat{\gamma}$ represents the background metric on which the deformed field theory is defined. The saddle-point metric $g_{\alpha}^{*}(\lambda)$ satisfies
$\frac{\delta}{\delta g_{\mu\nu}}\big{[}\hat{S}[g,\psi]+S^{(\lambda)}_{\mathrm{grav}}[g]\big{]}\big{|}_{g=g^{*}_{\alpha}}=0.$ (3)
@@
This framework incorporates the random geometric interpretation of two-dimensional $T\bar{T}$ deformation [14, 13]. The corresponding gravitational action is $S^{(\delta\lambda)}_{\mathrm{grav}}=(1/8\delta\lambda)\int d^{2}x\sqrt{\hat{\gamma}}\epsilon^{\mu\rho}\epsilon^{\nu\sigma}(g_{\mu\nu}-\hat{\gamma}_{\mu\nu})(g_{\rho\sigma}-\hat{\gamma}_{\rho\sigma})$, where $\gamma$ will eventually emerge as the background metric. By solving the EOM (3), the unique saddle is given by $g^{*}_{\mu\nu}=\hat{\gamma}_{\mu\nu}+2\delta\lambda\epsilon_{\mu\rho}\epsilon_{\nu\sigma}\hat{T}^{\rho\sigma}$. Substituting it into (4), the first-order $T\bar{T}$ deformed action is obtained. In this letter, the aforementioned framework is employed to investigate more general gravitational actions that involve the Ricci curvature and feature a dynamical metric.
@@ II.1 Effective field theory action on a deformed metric
Within the framework above, once the seed theory $\hat{S}$ and gravitational action $S^{(\lambda)}_{\mathrm{grav}}$ are specified, the deformed QFT action $S^{(\lambda)}_{\alpha}$ can, in principle, be obtained perturbatively from Eq. (4). However, determining the metric saddle point $g^{*}_{\alpha}(\lambda)$ nonperturbatively is generally intractable. Thus, it is useful to define an effective deformed action $S_{\mathrm{EFT}}^{(\lambda)}[\gamma^{(\lambda)},\psi]$ on a deformed metric $\gamma^{(\lambda)}$, which incorporates the saddle $g_{\alpha}^{*}$. The metric $\gamma^{(\lambda)}$ is taken as a local function of $\lambda$ and $g^{*}_{\alpha}$ (or, more generally, a functional of $g^{*}_{\alpha}$). Different choices of $S^{(\lambda)}_{\text{EFT}}$ and $\gamma^{(\lambda)}$ yield the same physical predictions when evaluated on $g^{*}_{\alpha}$.
A natural choice is $\gamma^{(\lambda)}\equiv g^{*}_{\alpha}$, in which case
$S_{\mathrm{EFT}}^{(\lambda)}[g_{\alpha}^{*},\psi]$ $=S_{\alpha}^{(\lambda)}[\hat{\gamma},\psi]$
@@
where $\mathcal{L}^{(\lambda)}_{\text{st}}$ is a local function of the seed theory stress tensor obtained by substituting the gravitational EOM into $S_{\mathrm{grav}}^{(\lambda)}$ and eliminating curvature terms.
Taking the total derivative of Eq. (4) with respect to $\lambda$ and using Eq. (3) yields
$\frac{\text{d}}{\text{d}\lambda}S_{\mathrm{EFT}}^{(\lambda)}[g_{\alpha}^{*},\psi]\equiv\frac{\text{d}}{\text{d}\lambda}S_{\alpha}^{(\lambda)}[\hat{\gamma},\psi]=\partial_{\lambda}S^{(\lambda)}_{\mathrm{grav}}[g^{*}_{\alpha}].$ (6)
This formulation provides a direct correspondence between a deformed QFT on the background $\hat{\gamma}$ and a classical gravitational theory, without relying on holography or extra dimensions. The deformation parameter $\lambda$ appears solely in the gravitational action, and the flow equation (6) relates the action’s variation to the underlying gravitational dynamics. For each stress tensor deformation, the corresponding gravitational action is given by relation (4).
III Linearized Einstein gravity
As a first application of the framework, the construction is implemented in Einstein gravity to make explicit how a stress-tensor deformation of the seed QFT reproduces linearized gravitational dynamics. The corresponding gravitational saddle and the non-local stress tensor deformation are explicitly derived.
The starting point is the seed theory coupled to the Einstein–Hilbert action,
$S=\hat{S}[g,\psi]+\frac{l^{2}}{2\lambda}\int d^{d}x\sqrt{g}\,(R-2\Lambda).$ (7)
The saddle-point equation for the metric yields the Einstein field equations with the field theory stress tensor as source,
$R^{*}_{\mu\nu}-\frac{1}{2}R^{*}g^{*}_{\mu\nu}+\Lambda g^{*}_{\mu\nu}=-\lambda l^{-2}\hat{T}_{\mu\nu}^{*},$ (8)
@@ III Linearized Einstein gravity
Consider the expansion $g=\hat{\gamma}+h$. By employing the standard second-order expansion of the Einstein–Hilbert action [22, 23], and including the gauge-fixing and ghost terms to handle diffeomorphism redundancy [24, 25], the expression for the saddle point $h^{*}$ in terms of the graviton Green’s function is obtained,
$h^{*}_{\mu\nu}(x)$ $=-\frac{2\lambda}{l^{2}}\int d^{d}y\sqrt{\hat{\gamma}(y)}G_{\mu\nu\rho\sigma}(x,y)\Big{(}\hat{T}^{\rho\sigma}(y)$
$\quad-\frac{1}{d-2}\hat{T}(y)\hat{\gamma}^{\rho\sigma}(y)\Big{)}+O(\lambda^{2}),$ (9)
with $G$ satisfying the equation $\hat{\gamma}^{\mu\rho}\hat{\gamma}^{\nu\sigma}\hat{\Box}-\hat{R}^{\mu\rho\nu\sigma}-\hat{R}^{\nu\rho\mu\sigma}]_{x}G_{\rho\sigma\alpha\beta}(x,y)=\delta^{(d)}(x\!-\!y)\delta^{\mu}_{\alpha}\delta^{\nu}_{\beta}/\!\sqrt{\!\hat{\gamma}}$. Solution (9) is obtained using the perturbation method, which is valid only when $\hat{\gamma}$ satisfies the vacuum Einstein equation. Besides, in the course of our calculation, some total derivative terms involving $h$ have been omitted. This can be realized by introducing suitable boundary terms into the action. The detailed derivations are shown in Supplemental Material A.
Substituting the saddle into the action yields the leading deformation of the QFT,
$S^{(\lambda)}$ $=\hat{S}+\frac{2\Lambda l^{2}\text{Vol}(\mathcal{M})}{\lambda(d-2)}+\frac{\lambda}{2l^{2}}\!\int\!d^{d}xd^{d}y\,\sqrt{\hat{\gamma}(x)\hat{\gamma}(y)}$
$\quad\times G_{\mu\nu\rho\sigma}(x,y)\,\hat{T}^{\mu\nu}(x)\Big{[}\hat{T}^{\rho\sigma}-\frac{1}{d-2}\hat{T}\hat{\gamma}^{\rho\sigma}\Big{]}(y)$
$\quad+O(\lambda^{2}).$ (10)
Similar non-local deformations were studied in [26]. For a flat background $\hat{\gamma}_{\mu\nu}=\eta_{\mu\nu}$ (with $\Lambda=0$), the first-order deformation simplifies to

### v2

# Geometric realization of stress-tensor deformed field theory

@@ Abstract
We present a semiclassical framework in which stress–tensor deformations of a quantum field theory (QFT) reorganize into a gravitational action evaluated at a metric saddle. The deformed partition function can be written as a gravitational path integral evaluated at the saddle, establishing a direct link between stress–tensor flows and gravitational dynamics. Two complementary routes arise: (i) from gravitational actions such as Einstein and Palatini, which map to stress–tensor deformations of a seed QFT; and (ii) from deformed QFTs such as generalized Nambu–Goto and $T\bar{T}$-like deformed models, which reconstruct the corresponding gravitational actions. Finally, in a free, massive scalar theory, we show that the one-loop effective action of the nonlocal deformation contains a local curvature term; its coefficient defines an induced Newton constant at a chosen renormalization scale, thereby demonstrating a bidirectional link between stress–tensor flows and classical gravity.
@@ I Introduction
Understanding whether and how spacetime geometry can be encoded in quantum field theory (QFT) remains a central problem in theoretical physics [44]. Thermodynamic arguments suggest that Einstein’s equations may reflect equilibrium conditions of quantum matter [24, 30], while holographic dualities such as AdS/CFT [27, 47] provide a framework for reconstructing spacetime from large-$N$ QFT data under specific boundary conditions, e.g.,[39]. Non-holographic proposals connecting geometric structures directly to QFT degrees of freedom [43] aim to broaden this perspective beyond fixed asymptotics.
Stress–tensor deformations provide a concrete probe of such connections. In classical gravity, $T_{\mu\nu}$ sources curvature, whereas in QFT, deformations built from $T_{\mu\nu}$ encode how geometry influences quantum matter. In two dimensions, the solvable $T\bar{T}$ deformation admits a precise geometric interpretation in terms of random geometry or topological gravity [48, 36, 13, 18, 41, 10]. Higher-dimensional generalizations have been explored [40, 9, 15], and recent works have constructed geometric realizations of stress–tensor flows without invoking dynamical gravity [6, 22]. These developments, however, primarily encode classical kinematical geometric features and do not yet incorporate propagating gravitational degrees of freedom.
This Letter formulates a calculable semiclassical correspondence in which the stress-tensor deformations reorganize a QFT into a gravitational action evaluated at a metric saddle, providing a geometric representation of stress–tensor deformed QFTs without claiming to derive gravity as a fundamental emergent phenomenon. Two complementary viewpoints arise. Starting from Einstein or Palatini gravity, one derives the corresponding stress–tensor deformation of the seed QFT. Conversely, starting from deformed theories such as generalized Nambu–Goto and $T\bar{T}$-like models, one reconstructs the associated gravitational action. Beyond the classical saddle, quantum corrections to a nonlocal stress–tensor deformation of massive scalar field theory produces a linearized Einstein gravitational action, showing that the classical metric saddle is the semiclassical limit of the same deformation and ensuring consistency between the classical and quantum levels. The correspondence is semiclassical and arises from a single QFT with a nonlocal stress–tensor deformation, distinguishing it from two-sector constructions [8], metric reformulations [15, 29], and flows derived from holographic renormalization [1].
@@ II Assumptions and Framework
The framework relates a stress–tensor–deformed QFT to a corresponding gravitational theory through the path integral, with the deformation controlled by a parameter $\lambda$. For simplicity, we mainly focus on the following Euclidean gravitational partition function is
@@
where $\hat{S}[g,\psi]$ is the seed theory action and $\psi$ denotes matter fields. The analysis presented in the first part of this Letter is conducted within a semiclassical framework. Evaluating the metric path integral at the saddle gives
@@
$=\int\mathcal{D}\psi\sum_{\hat{\gamma}}\!\sum_{\alpha}\mathcal{Z}^{(\lambda)}_{\mathrm{loops}}[g^{*}_{\alpha}]\,e^{-S^{(\lambda)}_{\alpha}[\hat{\gamma},\psi]},$ (2)
where $g^{*}_{\alpha}\to\hat{\gamma}$ as $\lambda\to 0$, and $\mathcal{Z}^{(\lambda)}_{\mathrm{loops}}[g^{*}]$ denotes the fluctuation (loop) contribution around each saddle $g_{\alpha}^{*}(\lambda)$ with $g_{\alpha}^{*}(\lambda)$ and $\hat{\gamma}$ satisfying
$\frac{\delta}{\delta g_{\mu\nu}}\big[\hat{S}[g,\psi]+S^{(\lambda)}_{\mathrm{grav}}[g]\big]\big|_{g=g^{*}_{\alpha}}$ $=0$
$\frac{\delta}{\delta g_{\mu\nu}}\big[\hat{S}[g,\psi]+S^{(\lambda=0)}_{\mathrm{grav}}[g]\big]\big|_{g=\hat{\gamma}}$ $=0.$ (3)
@@
This framework incorporates the random geometric interpretation of two-dimensional $T\bar{T}$ deformation [10, 41]. The corresponding gravitational action is $S^{(\delta\lambda)}_{\mathrm{grav}}=(1/8\delta\lambda)\int d^{2}x\sqrt{\hat{\gamma}}\epsilon^{\mu\rho}\epsilon^{\nu\sigma}(g_{\mu\nu}-\hat{\gamma}_{\mu\nu})(g_{\rho\sigma}-\hat{\gamma}_{\rho\sigma})$, where $\epsilon^{\mu\nu}$ is the Levi-Civita symbol, and $\hat{\gamma}$ will eventually emerge as the background metric. By solving the equations of motion (EOM) (3), the unique saddle is given by $g^{*}_{\mu\nu}=\hat{\gamma}_{\mu\nu}+2\delta\lambda\epsilon_{\mu\rho}\epsilon_{\nu\sigma}\hat{T}^{\rho\sigma}$, where $\hat{T}^{\rho\sigma}=\frac{-2}{\sqrt{\hat{\gamma}}}\frac{\delta\hat{S}}{\delta\hat{\gamma}_{\rho\sigma}}$. Substituting it into (4), the first-order $T\bar{T}$ deformed action is obtained. In this letter, the aforementioned framework is employed to investigate more general gravitational actions that involve the Ricci curvature and feature a dynamical metric. The construction provides a geometric representation of stress–tensor deformed QFTs, reproducing classical gravitational dynamics at the saddle.
@@ II.1 Effective field theory action on a deformed metric
Within the framework above, once the seed theory $\hat{S}$ and gravitational action $S^{(\lambda)}_{\mathrm{grav}}$ are specified, the deformed QFT action $S^{(\lambda)}_{\alpha}$ can, in principle, be obtained perturbatively from (4). However, determining the metric saddle point $g^{*}_{\alpha}(\lambda)$ nonperturbatively is generally intractable. Thus, it is useful to define an effective deformed action (EDA) $S_{\mathrm{EDA}}^{(\lambda)}$ on a deformed metric, which incorporates the saddle $g_{\alpha}^{*}$. The deformed metric is taken as a local function of $\lambda$ and $g^{*}_{\alpha}$ (more generally, a functional of $g^{*}_{\alpha}$).  [^1] Although this rewriting amounts to a saddle-level change of variables, it is operationally nontrivial: it recasts the intrinsically bilocal deformation into a covariant derivative expansion on the deformed geometry, yielding (i) a systematic perturbative control of subleading terms and (ii) a unified language for comparison with local stress–tensor flows of $T\bar{T}$-type. Once the Green’s-function prescription is fixed, scheme-dependent local counterterms are cleanly separated from the regulator-independent nonlocal kernel.
A natural choice is to take the metric saddle to be the deformed metric, in which case
$S_{\mathrm{EDA}}^{(\lambda)}[g_{\alpha}^{*},\psi]$ $=S_{\alpha}^{(\lambda)}[\hat{\gamma},\psi]$
@@
where $\mathcal{L}^{(\lambda)}_{\text{st}}$ is a local function of the seed theory stress-tensor obtained by substituting the gravitational EOM into $S_{\mathrm{grav}}^{(\lambda)}$ and eliminating terms associated with curvature tensors. Taking the total derivative of (4) with respect to $\lambda$ and using Eq. (3) yields
$\frac{\text{d}}{\text{d}\lambda}S_{\mathrm{EDA}}^{(\lambda)}[g_{\alpha}^{*},\psi]\equiv\frac{\text{d}}{\text{d}\lambda}S_{\alpha}^{(\lambda)}[\hat{\gamma},\psi]=\partial_{\lambda}S^{(\lambda)}_{\mathrm{grav}}[g^{*}_{\alpha}].$ (6)
This formulation provides a direct correspondence between a deformed QFT on the background $\hat{\gamma}$ and a classical gravitational theory. The deformation parameter $\lambda$ appears solely in the gravitational action, and the flow equation (6) relates the action’s variation to the underlying gravitational dynamics.
III Linearized Einstein gravity as stress tensor deformed theory
As a preliminary application, but not limited to, we apply the framework to linearized Einstein gravity. The metric saddle and the induced non-local stress-tensor deformation are derived explicitly. The starting point is the seed theory $\hat{S}[g,\psi]$ coupled to the Einstein–Hilbert action,
$S=\hat{S}[g,\psi]+\frac{l^{2}}{2\lambda}\int d^{d}x\sqrt{g}\,R.$ (7)
The saddle-point equation for the metric yields the Einstein field equations with the field theory stress-tensor as source,
$R^{*}_{\mu\nu}-\frac{1}{2}R^{*}g^{*}_{\mu\nu}=-\lambda l^{-2}\hat{T}_{\mu\nu}^{*},$ (8)
@@ III Linearized Einstein gravity as stress tensor deformed theory
Consider the expansion $g=\hat{\gamma}+\lambda h$. To keep the saddle-point fluctuation $h^{\ast}$ of order $O(\lambda^{0})$ and preserve the validity of the perturbative organization, we take the reference background $\hat{\gamma}_{\mu\nu}$ to satisfy the vacuum Einstein equations, as spelled out in SM A (A5). By employing the standard second-order expansion of the Einstein–Hilbert action [21, 2], and including the gauge-fixing and ghost terms to handle diffeomorphism redundancy with following standard procedure [14, 20] for Einstein gravities without cosmological constant  [^2] For dS and AdS gravities, the gauge-fixing can be found in the recent literature [28]., the solution for saddle point in terms of the graviton Green’s function is obtained  [^3] Solution of metric perturbation $h^{*}$ follows from a perturbative expansion around a vacuum Einstein background, which ensures the power counting $h^{\ast}_{\mu\nu}=O(\lambda^{0})$ so that higher-order metric terms do not affect the leading deformation. Total-derivative terms in $h$ are omitted, as they can be absorbed into boundary terms. Full derivations are provided in SM A (see in particular Eq. (A5)). $G_{\mu\nu\rho\sigma}$ is defined as the inverse of the gauge-fixed quadratic graviton operator obtained by expanding the gravitational action to second order around $\hat{\gamma}_{\mu\nu}$. Ghosts contribute only to the one-loop determinant and do not affect the tree-level saddle equation; see SM A., $h^{*}_{\mu\nu}=-2l^{-2}\int d^{d}y\sqrt{\hat{\gamma}}G_{\mu\nu\rho\sigma}(\hat{T}^{\rho\sigma}-\frac{1}{d-2}\hat{T}\hat{\gamma}^{\rho\sigma})+O(\lambda)$, with $G$ satisfying the equation $\hat{\gamma}^{\mu\rho}\hat{\gamma}^{\nu\sigma}\hat{\Box}-\hat{R}^{\mu\rho\nu\sigma}-\hat{R}^{\nu\rho\mu\sigma}]_{x}G_{\rho\sigma\alpha\beta}(x,y)=\delta^{(d)}(x\!-\!y)\delta^{\mu}_{\alpha}\delta^{\nu}_{\beta}/\!\sqrt{\!\hat{\gamma}}$  [^4] For generic curved background, one can define the reference background $\hat{\gamma}$ satisfies (3) and the Green’s function can be defined in a classical way which highly depends on the boundary conditions. For simplicity, here we consider the Green’s function on a manifold without boundary, which exhibits exponential decay as the separation between the two points tends to infinity.. Substituting the saddle into the action yields the leading deformation of the QFT,
$S^{(\lambda)}$ $=\hat{S}+\frac{\lambda}{2l^{2}}\!\int\!d^{d}xd^{d}y\,\sqrt{\hat{\gamma}(x)\hat{\gamma}(y)}G_{\mu\nu\rho\sigma}(x,y)$
$\quad\times\,\hat{T}^{\mu\nu}(x)\Big[\hat{T}^{\rho\sigma}-\frac{1}{d-2}\hat{T}\hat{\gamma}^{\rho\sigma}\Big](y)+O(\lambda^{2}).$ (9)
A standard de Donder gauge-fixing of the metric fluctuation (including the associated ghost term) is employed (SM A (A7)). $G_{\mu\nu\rho\sigma}(x,y)$ is the Green’s function of the resulting gauge-fixed spin-2 operator on the reference background $\hat{\gamma}_{\mu\nu}$, with boundary conditions chosen according to the manifold under consideration. At the order relevant for (III), $G_{\mu\nu\rho\sigma}$ enters only through its first part when contracted with the conserved source $\hat{T}_{\mu\nu}$; gauge-dependent longitudinal components therefore do not affect the deformation. Similar non-local deformations were studied in [25]. For a flat background $\hat{\gamma}_{\mu\nu}=\eta_{\mu\nu}$, the first-order deformation simplifies to